# Differentiation, the Exception Not the Rule - Evidence for Full Miscibility in Sub-Neptune Interiors

### v1

@@ Abstract
We investigate the consequences of non-ideal mixing between silicate, iron metal, and hydrogen for the structures of the cores of sub-Neptunes with implications for super-Earths, warm Neptunes, and ice giants. A method of extrapolating what we know about the miscibility in the three bounding binary systems MgSiO3-H2, MgSiO3-Fe, and Fe-H2 to the ternary composition space is used to deduce the phase equilibria of this system at relevant temperature and pressure conditions. We find that while separate silicate and metal phases can exist at shallow depths, the phases become entirely miscible deeper in the cores, thus altering the density structure of the cores. The assumption that the interiors of large rocky planets, either with extant magma oceans beneath H2-rich envelopes, or evolved from such bodies, are composed of a differentiated metal core overlain by a silicate mantle is inconsistent with our understanding of the phase equilibria of these bodies.
@@ 1 Introduction
Based on the Kepler survey, sub-Neptunes are the most abundant planet type in the Galaxy (e.g., Fressin et al., 2013). These bodies have bulk densities, $\rho$, similar to about 1 to 3 g cm-3 compared with Earth’s bulk density of $5.5$ g cm-3. Studies of the atmospheres of sub-Neptunes are multiplying rapidly in the age of the James Webb Space Telescope (JWST). However, our understanding of their cores, meaning everything beneath the outer envelope, is necessarily limited. In some senses, we are in a period of discovery for sub-Neptunes that parallels our understanding of the core of the Sun in the 1920s (Payne-Gaposchkin, 1925). We need judicious use of physical chemistry to infer what is beneath the outer veils of these copious planets.
@@
Alternatively, sub-Neptunes may have rock-like, molten or partially molten interiors beneath hydrogen-rich primary atmospheres (Owen & Wu, 2013; Rogers et al., 2023). The H2-rich envelopes in this interpretation must comprise percent levels of the planet masses in order to match their bulk densities (e.g. Bean et al., 2021). For example, assuming a density for H2 of 0.1 g cm-3, appropriate for the high pressures and temperatures at the base of the envelopes (e.g., Chabrier et al., 2019), a compressed core density similar to that for the whole Earth, and a median sub-Neptune density of 2 g cm-3, the planet would have an H2 envelope comprising $2.7$ of its mass.
@@
The bulk densities of super-Earths indicate that they are rocky planets with broadly Earth-like bulk densities and negligible atmospheres in terms of mass. A strong case can be made that super-Earths form when some sub-Neptunes lose their dense hydrogen-rich envelopes (Rogers et al., 2024), leaving behind rocky cores. This leads to the well-known radius gap that separates these two classes of planets (Fulton et al., 2017; Owen & Wu, 2013; Gupta & Schlichting, 2019): super-Earths have radii between $>1$ and $1.7R_{\oplus}$, while sub-Neptunes range from about $1.9$ to $4.0R_{\oplus}$. In this view, super-Earth interiors may be similar to those of sub-Neptunes (Bean et al., 2021). An important caveat is the prospect for degassing of H2 from the molten interior of a sub-Neptune as it loses its envelope and cools inexorably below some threshold final equilibration temperature (e.g., Rogers et al., 2025). While super-Earths appear to evolve from sub-Neptunes, it is not yet clear that they retain all of the properties of sub-Neptune cores.
@@
Our paper is structured as follows. In section 2 we discuss bulk chemical constraints on iron-rich cores for exoplanets in general. In section 3 we describe the binary mixing relationships that are the basis for our analysis, and section 4 presents the ternary extension of these mixing models. The resulting ternary phase diagram is presented in section 5 and implications for the interior structures of sub-Neptunes are discussed in section 6. Conclusions are presented in the final section.
@@ 3.2 MgSiO3-H2 system
Here, $L_{\mathrm{CB}}$ and $L_{\mathrm{BC}}$ are the binary interaction parameters where the logic for the labels CB and BC will be apparent when we apply these to our ternary model, and $x$ is the mole fraction of H2 along the binary join. In this formulation, the excess, or non-ideal, molar enthalpy of mixing is $\Delta\hat{H}_{\mathrm{excess}}=\left(L_{\mathrm{CB}}x+L_{\mathrm{BC}}(1-x)\right)x(1-x)$. The temperature and pressure dependence for the non-ideal mixing is embodied by excess entropy and volume parameters, represented by $\tau$ and $\pi$ that are $\Delta\hat{H}_{\mathrm{excess}}/\hat{S}_{\mathrm{excess}}$ and $\Delta\hat{H}_{\mathrm{excess}}/\Delta\hat{V}_{\mathrm{excess}}$, respectively (Stixrude & Gilmore, 2025). The values for the subregular solution fit to the DFT simulations are: $L_{\mathrm{CB}}=786000$ J/mol, $L_{\mathrm{BC}}=-6260$ J/mol, $\tau=4670$ K, and $\pi=-35$ GPa.
@@
We provide an example of the basis for this non-ideal mixing model using two illustrative DFT-MD simulations that bracket the position of the binodal (Figures 3 and 4). The initial condition for these simulations is composed of two distinct, adjacent phases, MgSiO3 melt and H2 gas, at the same temperature and about the same pressure (within about 5%). The hydrogen in each case comprises 4% by weight of the system, corresponding to a composition near the peak of the binodal. After suitably long simulation times of $>13$ ps, results show that the silicate-H2 system is completely miscible at 4,000 K and $3.5$ GPa (Figure 3), but not at 3,000 K and $2.5$ GPa (Figure 4 ), where the different pressures were used to preserve number of atoms and volumes for the two different temperatures. This shows that the implied $T$ and $P$ for the binodal for this bulk composition are indeed near those expected for sub-Neptunes with of order a few weight per cent H2.
@@ 3.3 MgSiO3-Fe system
To our knowledge there are no published analogous mixing calculations for MgSiO3-Fe. Nonetheless, it is reasonable to infer that there is non-ideal mixing along this join, as in the case of MgO-Fe, and that there is a well-defined binodal that closes at sufficiently high temperatures and pressures. In order to test whether the regular solution model for MgO-Fe mixing can be used as a surrogate for MgSiO3-Fe mixing at high $T$ and $P$, we performed DFT molecular dynamics simulations of MgSiO3-Fe mixing. Results show that at conditions above the MgO-Fe binodal, MgSiO3 and Fe are completely miscibile (Figure 5).
@@ 3.4 Fe-H2 system
Constraints on the thermodynamics of mixing between Fe and H2 at relevant $T$ and $P$ come from experiments and ab initio modeling. At relatively low temperatures and pressures (e.g., 2000 K, 0.1 GPa), iron-rich melt coexists with H2 gas (San-Martin & Manchester, 1990). There is, however, both experimental and modeling evidence that the solubility of H2 in liquid iron rises sharply at high $T$ and $P$. Tagawa et al. (2021) performed experiments measuring the distribution coefficient for H in molten Fe metal relative to H in molten silicate, $D_{\mathrm{H}}=[H]_{\mathrm{metal}}/[H]_{\mathrm{silicate}}$, from 3100 to 4600 K and 30 to 60 GPa. They found that while $D_{\mathrm{H}}$ is near unity at $T<3000$ K, it increases to 10 at 3000 K and 30 GPa and to 30 at 4000 K and 60 GPa. Similarly, Yuan & Steinle-Neumann (2020) performed ab initio modeling showing that $D_{\mathrm{H}}$ increases from near unity at 0 GPa to about 100 at 40 GPa. These observations all point to an enhancement in the solubility of H2 in molten Fe metal with pressure. The requirement for progressive closure of the miscibility gap between H2 and liquid iron with pressure, and the relative distribution of H between molten silicate and iron (increasing $D_{\mathrm{H}}$) permit a rough calibration of the Fe-H2 mixing relations. We find that these relationships are reproduced using a subregular solution model where the interaction parameters are pressure dependent. The model is
@@
$RT\left(x\ln x+(1-x)\ln(1-x)\right).$
where $L_{\mathrm{AB}}=115000-9500P(\mathrm{GPa})$ J/mol, $L_{\mathrm{BA}}=17000-9500P(\mathrm{GPa})$ J/mol, and $x$ is the mole fraction of H2 on the binary join. While the precise values used here are certainly not unique, the general shape of the free energy of mixing surface is required by the observations described above. For example, with a weaker pressure dependence, the enhancement in the siderophile nature of H with pressure is not captured.
@@ 4 Ternary mixing model
The basis for our analysis is the topology of the Gibbs free energy of mixing surface in the ternary system MgSiO3-Fe-H2. We will make use of the mixing models for the three binary systems MgSiO3-H2, Fe-H2, and MgSiO3-Fe and extrapolate these bounding binary mixing models into the ternary system. There is a long history of performing such extrapolations with success with percent levels of accuracy (Wohl, 1946, 1953; Muggianu et al., 1975; Jacob & Fitzner, 1977; Lee & Kim, 2001). We adopt the Muggianu-Jacob method of projecting along shortest distances from each of the binaries to any given point in the ternary system (Muggianu et al., 1975; Jacob & Fitzner, 1977).
@@
$x_{\mathrm{A}}x_{\mathrm{B}}\left(L_{\mathrm{AB}}\cdot\frac{1}{2}(1+x_{\mathrm{B}}-x_{\mathrm{A}})+L_{\mathrm{BA}}\cdot\frac{1}{2}(1+x_{\mathrm{A}}-x_{\mathrm{B}})\right)$
$+x_{\mathrm{B}}x_{\mathrm{C}}\left(L_{\mathrm{BC}}\cdot\frac{1}{2}(1+x_{\mathrm{C}}-x_{\mathrm{B}})+L_{\mathrm{CB}}\cdot\frac{1}{2}(1+x_{\mathrm{B}}-x_{\mathrm{C}})\right)$
@@
$+x_{\mathrm{C}}x_{\mathrm{A}}\left(L_{\mathrm{CA}}\cdot\frac{1}{2}(1+x_{\mathrm{A}}-x_{\mathrm{C}})+L_{\mathrm{AC}}\cdot\frac{1}{2}(1+x_{\mathrm{C}}-x_{\mathrm{A}})\right)$
@@
where $L_{ij}$ are the interaction parameters for species $i$ and $j$, and $\tau$ and $\pi$ parameterize the temperature and pressure dependencies of the non-ideal mixing parameters for the MgSiO3-H2 subregular solution model (Stixrude & Gilmore, 2025). Expressions of the form $\frac{1}{2}(1+x_{i}-x_{j})$ are the normal projections of the mole fractions of the independent component $i$ onto the binary $i-j$ join.
@@ 4.2 Binodals
The binodal pairs along the coexistence curves in ternary composition space at specified $T$ and $P$ are defined by applying the criteria for common tangents, satisfying the requirement that $\mu_{i}^{\alpha}=\mu_{j}^{\beta}$ where $\mu_{i}^{\alpha}$ is the chemical potential for species $i$ in phase $\alpha$ and $\mu_{i}^{\beta}$ is the chemical potential for that same component in phase $\beta$. The chemical potentials are defined as
$\mu_{i}^{\alpha}=\frac{\partial\hat{G}_{\mathrm{mix}}}{\partial x_{\mathrm{i}}^{\alpha}}$ (7)
for phase $\alpha$ and similarly for phase $\beta$.
@@ 5.1 Phase changes with P and T
Figure 6: Gibbs free energy of mixing surface, $\hat{G}_{\rm mix}$, in the ternary system MgSiO3-Fe-H2 for $T=3569$ K and $P=3.6$ GPa. The heavy black lines show the spinodals denoting the regions of spontaneous decomposition into two phases. Points comprising the binodals are shown in blue connected by tie lines. The blue binodal points define the coexistence curves. Between the binodals and spinodals lies the region of metastability for two phases.
@@
Figure 6 shows the resulting Gibbs free energy surface, $\hat{G}_{\rm mix}$, in the ternary system MgSiO3-Fe-H2 for $T=3569$ K and $P=3.6$ GPa. The ternary coordinates are the mole fractions of the three components. The heavy black lines show the spinodals. Interior to the spinodals the curvature of $\hat{G}_{\rm mix}$ is negative, and any infinitesimal perturbation lowers the free energy. Here, phase separation is spontaneous. Each of the coexisting stable phases lies at a point in the phase diagram where the curvature in $\hat{G}_{\rm mix}$ is positive and the points have the same chemical potentials (i.e., equal $\partial\hat{G}_{\mathrm{mix}}/\partial x_{\mathrm{i}}$). These points, the binodals, are shown as blue points connected by tie lines. They define the coexistence curves for two phases in the phase diagrams. Between the binodals and spinodals lies the region of metastability for two phases.
@@
An example of an isothermal, isobaric phase diagram in ternary space is shown in Figure 7. The diagram is a projection of the free energy surface in the system MgSiO3-Fe-H2 at $T=4128$ K and $P=12.5$ GPa. Ternary coordinates are mole fractions of the three components. Regions of 1-phase stability outside of both the spinodal and binodal curves are shown in pale grey. Regions inside of the spinodals are where two liquid phases must occur and are shown in darker grey. The binodals are shown by the blue curves and the compositions of coexisting two stable phases are shown by the tie lines connecting the binodes. The regions between the spinodals and binodals are where two phases might persist metastably and are shown in sky blue. The $\oplus$ symbol represents the bulk composition of Earth assuming $0.5$ weight percent H2 in its metal core (Young et al., 2023). The [Uncaptioned image] symbol represents a fiducial sub-Neptune composition for a bulk H2 concentration of 2% by mass and the remainder of the planet composed of MgSiO3 silicate and Fe in 2:1 mass proportions. The projected position of Neptune in this space, [Uncaptioned image], is also shown, assuming the bulk H2 concentration is 20% by weight (approximately $0.9$ in mole fraction).
@@
Figure 7: A ternary phase diagram for the system MgSiO3-Fe-H2 at $T=5318$ K and $P=10.5$ GPa. Regions of 1-phase stability that lie outside of both the spinodal and binodal curves are shown in pale grey. Regions inside of the spinodals where two liquid phases are ensured are shown in darker grey. The binodals are shown by the blue curves and coexisting phase compositions for a given bulk composition are shown by the tie lines connecting individual binodes. Regions of two-phase metastability are between the spinodals and binodals and are shown in the sky blue color. Reference planet bulk compositions projected into this ternary space, in mole fractions, are shown for Earth ($\oplus$), our fiducial sub-Neptune (Refer to caption), and Neptune (Refer to caption), with weight fractions of H2 being 0.17%, 2%, and 20%, respectively.
@@
We constructed a fiducial sub-Neptune planet with a total mass of $6M_{\oplus}$ and a bulk H2 concentration of 2% by mass. We use the equations of state (EoS) for MgSiO3, Fe, and H2 (see Appendix B for material properties) where the surface that separates the molten core from the outer H2-rich envelope is the binodal in the system MgSiO3-H2 as obtained by Stixrude & Gilmore (2025). The surface of the supercritical magma ocean is therefore as described by Young et al. (2024). This surface moves inward as the planet cools, resulting in a lower temperature and higher pressure for the phase boundary; the conditions at the binodal separating the magma ocean from the envelope are tied to the overall thermal state of the planet.
@@
Figure 8: Pressure as a function of radial position along an isentrope in the interior of our fiducial sub-Neptune with a fully miscible core. The envelope-core interface is at $R=2R_{\oplus}$. The planet total mass is $6M_{\oplus}$ and has a bulk H2 concentration of 2% by mass. The positions of some of the panels shown in Figure 9 are shown for reference.
For these calculations, we prescribe the mass of the planet, the total H2 abundance, the equilibrium temperature for the atmosphere, and the thermal state of the planet as expressed by the pressure of the binodal defining the outer boundary of the magma ocean, $P_{\rm S}$. The pressure of the binodal is a manifestation of the total energy of the planet. Solutions are obtained by iterating between the core and atmosphere structures as follows. As a starting point for the iteration, we assume all of the H2 resides in the core and calculate the surface temperature of the magma ocean, i.e. the binodal temperature where one phase becomes unstable relative to two stable phases, melt and envelope, at the prescribed concentration of H2. This temperature is the reduced temperature for the adiabat in the core (i.e., the surface $T_{\rm S}$ for given $P_{\rm S}$). The shape of the binodal determines the composition of the atmosphere in equilibrium with the melt at $T_{\rm S}$ and $P_{\rm S}$ (Figure 2). From the pressure at the base of the atmosphere (the pressure of the binodal) and its composition, we calculate an atmosphere model, including its mass. The lower boundary conditions for the atmosphere are set by the magma-envelope binodal, and the upper boundary is set by the equilibrium temperature. We then subtract the mass of hydrogen in this atmosphere from the core, and revise the core calculation, repeating for several iterations until the planet structure converges.

### v4

@@ Abstract
We investigate the consequences of non-ideal mixing between silicate, iron metal, and hydrogen for the structures of the cores of sub-Neptunes with implications for super-Earths and perhaps other planets formed by mixtures of silicate melt, iron metal, and significant mass fractions of hydrogen. A method of extrapolating what we know about the miscibility in the three bounding binary systems MgSiO3-H2, MgSiO3-Fe, and Fe-H2 to the ternary composition space is used to deduce the phase equilibria of this system at relevant temperature and pressure conditions. We find that while separate silicate and metal phases can exist at shallow depths, the phases become entirely miscible deeper in the cores, thus altering the density structure of the cores. The assumption that the interiors of large rocky planets, either with extant magma oceans beneath H2-rich envelopes, or evolved from such bodies, are composed of a differentiated metal core overlain by a silicate mantle is inconsistent with our understanding of the phase equilibria of these bodies.
@@ 1 Introduction
Based on the Kepler survey, sub-Neptunes are the most abundant planet type in the Galaxy (e.g., Fressin et al., 2013; Fulton et al., 2017). These bodies have bulk densities, $\rho$, similar to about 1 to 3 g cm-3 compared with Earth’s bulk density of $5.5$ g cm-3. Studies of the atmospheres of sub-Neptunes are multiplying rapidly in the age of the James Webb Space Telescope (JWST). However, our understanding of their cores, meaning everything beneath the outer envelope, is necessarily limited. In some senses, we are in a period of discovery for sub-Neptunes that parallels our understanding of the core of the Sun in the 1920s (Payne-Gaposchkin, 1925). We need judicious use of physical chemistry to infer what is beneath the outer veils of these copious planets.
@@
Alternatively, sub-Neptunes may have rock-like, molten or partially molten interiors beneath hydrogen-rich primary atmospheres (Chachan & Stevenson, 2018). The H2-rich envelopes in this interpretation must comprise percent levels of the planet masses in order to match their bulk densities (e.g. Bean et al., 2021). For example, assuming a density for H2 of 0.1 g cm-3, appropriate for the high pressures and temperatures at the base of the envelopes (e.g., Chabrier et al., 2019), a compressed core density similar to that for the whole Earth, and a median sub-Neptune density of 2 g cm-3, the planet would have an H2 envelope comprising $2.7$ % of its mass.
@@
The bulk densities of super-Earths indicate that they are rocky planets with broadly Earth-like bulk densities and negligible atmospheres in terms of mass. A strong case can be made that super-Earths form when some sub-Neptunes lose their dense hydrogen-rich envelopes (Rogers et al., 2024), leaving behind rocky cores. This leads to the well-known radius gap that separates these two classes of planets (Fulton et al., 2017; Owen & Wu, 2013; Gupta & Schlichting, 2019): super-Earths have radii between >1>1 and $1.7R_{\oplus}$, while sub-Neptunes range from about $1.9$ to $4.0R_{\oplus}$. In this view, super-Earth interiors may be similar to those of sub-Neptunes (Bean et al., 2021). An important caveat is the prospect for degassing of H2 from the molten interior of a sub-Neptune as it loses its envelope and cools inexorably below some threshold final equilibration temperature (e.g., Rogers et al., 2025). While super-Earths appear to evolve from sub-Neptunes, it is not yet clear that they retain all of the properties of sub-Neptune cores.
@@
Our paper is structured as follows. In section 2 we discuss bulk chemical constraints on iron-rich cores for exoplanets in general. In section 3 we describe the binary mixing relationships that are the basis for our analysis. In this section we present existing and newly-derived binary Gibbs free energy surfaces, and we present three new ab initio - molecular dynamics simulations in support of these free energy surfaces. In section 4 we present our new ternary Gibbs free energy surface based on the binary systems, and a convex hull approach for calculating the stable phase equilibria in the ternary system. The resulting ternary phase diagram is presented in section 5 and implications for the interior structures of sub-Neptunes are discussed in section 6. Conclusions are presented in the final section.
@@ 3.2 MgSiO3-H2 system
Here, $L_{\mathrm{CB}}$ and $L_{\mathrm{BC}}$ are the binary interaction parameters where the logic for the labels CB and BC will be apparent when we apply these to our ternary model, and $x$ is the mole fraction of H2 along the binary join. In this formulation, the excess, or non-ideal, molar enthalpy of mixing is $\Delta\hat{H}_{\mathrm{excess}}=\left(L_{\mathrm{CB}}x+L_{\mathrm{BC}}(1-x)\right)x(1-x)$. The temperature and pressure dependence for the non-ideal mixing is embodied by excess entropy and volume parameters, represented by $\tau$ and $\pi$ that are $\Delta\hat{H}_{\mathrm{excess}}/\hat{S}_{\mathrm{excess}}$ and $\Delta\hat{H}_{\mathrm{excess}}/\Delta\hat{V}_{\mathrm{excess}}$, respectively (Stixrude & Gilmore, 2025). The values for the subregular solution fit to the DFT simulations are: $L_{\mathrm{CB}}=622000$ J/mol, $L_{\mathrm{BC}}=-4950$ J/mol, $\tau=4800$ K, and $\pi=-35$ GPa.
@@
Here we present two new DFT-MD simulations that bracket the position of the binodal (Figures 3 and 4) as a check on the veracity of the model and an illustration of the two-phase mixing calculations. Methods for the DFT-MD calculations are described in Appendix A. The initial condition for these simulations is composed of two distinct, adjacent phases, MgSiO3 melt and H2 gas, at the same temperature and about the same pressure (within about 5%). The hydrogen in each case comprises 4% by weight of the system, corresponding to a composition near the peak of the binodal. After suitably long simulation times of >13>13 ps, results show that the silicate-H2 system is completely miscible at 4,000 K and $3.5$ GPa (Figure 3), but not at 3,000 K and $2.5$ GPa (Figure 4 ), where the different pressures were used to preserve number of atoms and volumes for the two different temperatures. This shows that the implied $T$ and $P$ for the binodal for this bulk composition are indeed near those expected for sub-Neptunes with of order a few weight per cent H2.
@@ 3.3 MgSiO3-Fe system
To our knowledge there are no published analogous mixing calculations for MgSiO3-Fe. Nonetheless, it is reasonable to infer that there is non-ideal mixing along this join, as in the case of MgO-Fe, and that there is a well-defined binodal that closes at sufficiently high temperatures and pressures.
We performed DFT-MD simulations of the MgSiO3-Fe system in order to test whether the regular solution model for MgO-Fe mixing can be used as a surrogate for the MgSiO3-Fe system at high $T$ and $P$. Details of the calculations are presented in Appendix A. Results show that at conditions above the MgO-Fe binodal, MgSiO3 and Fe are indeed completely miscibile (Figure 5).
@@ 3.4 Fe-H2 system
Constraints on the thermodynamics of mixing between Fe and H2 at relevant $T$ and $P$ come from experiments and ab initio modeling. At relatively low temperatures and pressures (e.g., 2000 K, 0.1 GPa), iron-rich melt coexists with H2 gas (San-Martin & Manchester, 1990). There is, however, both experimental and modeling evidence that the solubility of H2 in liquid iron rises sharply at high $T$ and $P$. Tagawa et al. (2021) performed experiments measuring the distribution coefficient for H in molten Fe metal relative to H in molten silicate, $D_{\mathrm{H}}=[H]_{\mathrm{metal}}/[H]_{\mathrm{silicate}}$, from 3100 to 4600 K and 30 to 60 GPa. They found that while $D_{\mathrm{H}}$ is near unity at $T<3000$ K, it increases to 10 at 3000 K and 30 GPa and to 30 at 4000 K and 60 GPa. Similarly, Yuan & Steinle-Neumann (2020) performed ab initio modeling showing that $D_{\mathrm{H}}$ increases from near unity at 0 GPa to about 100 at 40 GPa. These observations all point to an enhancement in the solubility of H2 in molten Fe metal with pressures greater than about 10 GPa. The requirement for progressive closure of the miscibility gap between H2 and liquid iron with pressure Fukai (1992), and the relative distribution of H between molten silicate and iron (increasing $D_{\mathrm{H}}$) permit a rough calibration of the Fe-H2 mixing relations.
An additional constraint comes from the parameterization of the solubility of H in molten Fe at lower temperatures and pressures. The compilations by Zhang (2007) and Jiang et al. (2011) yield the expression
[…]
We make use of this Equation as a constraint on the Gibbs free energy surface for mixing of molten Fe and H2 by matching the melt limb of the binodal to the solubility expressed by Equation 3 at conditions appropriate for our fiducial sub-Neptune magma ocean-H2-rich envelop interface (the minimum required extrapolation of $T$ and $P$).
We find that all of these relationships are reproduced using a subregular solution model where the interaction parameters are pressure dependent. The model is
@@
$RT\left(x\ln x+(1-x)\ln(1-x)\right),$
where $L_{\mathrm{AB}}=138000-9500P(\mathrm{GPa})$ J/mol, $L_{\mathrm{BA}}=17000-9500P(\mathrm{GPa})$ J/mol, and $x$ is the mole fraction of H2 on the binary join. The resulting isobaric phase diagram at 4 GPa is shown in Figure 6. While the precise values used here are certainly not unique, the general shape of the free energy of mixing surface is required by the observations described above.
Figure 6: Isobaric phase diagram for the Fe–H2 binary system based on Equation 4. The extrapolation of the low-$P$ and low-$T$ hydrogen solubility in liquid Fe metal is shown for comparison (Zhang, 2007; Jiang et al., 2011). Errors for the latter are derived by standard propogation of uncertainties.
@@ 4 Ternary mixing model
We calculate the topology of the Gibbs free energy of mixing in the ternary system MgSiO3-Fe-H2 as the basis for our analysis. We make use of the mixing models for the three binary systems MgSiO3-H2, Fe-H2, and MgSiO3-Fe and extrapolate these bounding binary mixing models into the ternary system. There is a long history of performing such extrapolations with success with percent levels of accuracy (Wohl, 1946, 1953; Muggianu et al., 1975; Jacob & Fitzner, 1977; Lee & Kim, 2001). We adopt the Muggianu-Jacob method of projecting along shortest distances from each of the binaries to any given point in the ternary system (Muggianu et al., 1975; Jacob & Fitzner, 1977).
@@
$x_{\mathrm{A}}x_{\mathrm{B}}\left(L_{\mathrm{AB}}\frac{1}{2}(1+x_{\mathrm{B}}-x_{\mathrm{A}})+L_{\mathrm{BA}}\frac{1}{2}(1+x_{\mathrm{A}}-x_{\mathrm{B}})\right)$
$+x_{\mathrm{B}}x_{\mathrm{C}}\left(L_{\mathrm{BC}}\frac{1}{2}(1+x_{\mathrm{C}}-x_{\mathrm{B}})+L_{\mathrm{CB}}\frac{1}{2}(1+x_{\mathrm{B}}-x_{\mathrm{C}})\right)$
@@
$+x_{\mathrm{C}}x_{\mathrm{A}}\left(L_{\mathrm{CA}}\frac{1}{2}(1+x_{\mathrm{A}}-x_{\mathrm{C}})+L_{\mathrm{AC}}\frac{1}{2}(1+x_{\mathrm{C}}-x_{\mathrm{A}})\right)$
@@
where $L_{ij}$ are the interaction parameters for species $i$ and $j$, and $\tau$ and $\pi$ parameterize the temperature and pressure dependencies of the non-ideal mixing parameters for the MgSiO3-H2 subregular solution model (Stixrude & Gilmore, 2025). Expressions of the form $\frac{1}{2}(1+x_{i}-x_{j})$ are the normal projections of the mole fractions of the independent component $i$ onto the binary iji-j join.
@@ 4.2 Binodals
The binodal pairs along the coexistence curves in ternary composition space at specified $T$ and $P$ are defined by applying the criteria for common tangents, satisfying the requirement that $\mu_{i}^{\alpha}=\mu_{i}^{\beta}$ where $\mu_{i}^{\alpha}$ is the chemical potential for species $i$ in phase $\alpha$ and $\mu_{i}^{\beta}$ is the chemical potential for that same component in phase $\beta$. The mixing chemical potentials are related by the first derivative of the free energy of mixing
$\mu_{i,{\rm mix}}^{\alpha}-\mu_{j,{\rm mix}}^{\alpha}=\frac{\partial\hat{G}_{\mathrm{mix}}}{\partial x_{\mathrm{i}}^{\alpha}}$ (8)
for independently variable component $i$ and dependent component $j$ in phase $\alpha$ and similarly for phase $\beta$.
@@ 5.1 Phase changes with P and T
Figure 7: Gibbs free energy of mixing surface, $\hat{G}_{\rm mix}$, in the ternary system MgSiO3-Fe-H2 for $T=3500$ K and $P=4$ GPa. The heavy black lines show the spinodals denoting the regions of spontaneous decomposition into two phases. Points comprising the binodals are shown in blue connected by tie lines. The blue binodal points define the coexistence curves. Between the binodals and spinodals lies the region of metastability for two phases.
@@
Figure 7 shows the resulting Gibbs free energy surface, $\hat{G}_{\rm mix}$, in the ternary system MgSiO3-Fe-H2 for $T=3500$ K and $P=4$ GPa where conditions were chosen for illustration purposes only. The ternary coordinates are the mole fractions of the three components. The heavy black lines show the spinodals. Interior to the spinodals the curvature of $\hat{G}_{\rm mix}$ is negative, and any infinitesimal perturbation lowers the free energy. Here, phase separation is spontaneous. Each of the coexisting stable phases lies at a point in the phase diagram where the curvature in $\hat{G}_{\rm mix}$ is positive and the points have the same chemical potentials (i.e., equal $\partial\hat{G}_{\mathrm{mix}}/\partial x_{\mathrm{i}}$). These points, the binodals, are shown as blue points connected by tie lines. They define the coexistence curves for two phases in the phase diagrams. Between the binodals and spinodals lies the region of metastability for two phases.
@@
An example of an isothermal, isobaric phase diagram in ternary space is shown in Figure 8. The diagram is a projection of the free energy surface in the system MgSiO3-Fe-H2 at $T=5318$ K and $P=13.5$ GPa, where the $T$ and $P$ were again chosen for illustration purposes. Ternary coordinates are mole fractions of the three components. Regions of 1-phase stability outside of both the spinodal and binodal curves are shown in pale grey. Regions inside of the spinodals are where two liquid phases must occur and are shown in darker grey. The binodals are shown by the blue curves and the compositions of coexisting two stable phases are shown by the tie lines connecting the binodes. The regions between the spinodals and binodals are where two phases might persist metastably and are shown in sky blue. The $\oplus$ symbol represents the bulk composition of Earth assuming $0.5$ weight percent H2 in its metal core (Young et al., 2023). The [Uncaptioned image] symbol represents a fiducial sub-Neptune composition for a bulk H2 concentration of 2% by mass and the remainder of the planet composed of MgSiO3 silicate and Fe in 2:1 mass proportions. The projected position of Neptune in this space, [Uncaptioned image], is also shown, assuming the bulk H2 concentration is 20% by weight (approximately $0.9$ in mole fraction). Where the hydrogen-rich envelopes do not dominate the mass of H2, the bulk planet and bulk interior hydrogen concentrations are similar, although not identical.
@@
Figure 8: A ternary phase diagram for the system MgSiO3-Fe-H2 at $T=5318$ K and $P=13.5$ GPa. Regions of 1-phase stability that lie outside of both the spinodal and binodal curves are shown in pale grey. Regions inside of the spinodals where two liquid phases are ensured are shown in darker grey. The binodals are shown by the blue curves and coexisting phase compositions for a given bulk composition are shown by the tie lines connecting individual binodes. Regions of two-phase metastability are between the spinodals and binodals and are shown in the sky blue color. Reference planet bulk compositions projected into this ternary space, in mole fractions, are shown for Earth ($\oplus$), our fiducial sub-Neptune (Refer to caption), and Neptune (Refer to caption), with weight fractions of H2 being 0.17%, 2%, and 20%, respectively.
@@
We constructed a new fiducial model sub-Neptune planet with a total mass of $6M_{\oplus}$ and a bulk H2 concentration of 2% by mass based on the equations of state (EoS) for MgSiO3, Fe, and H2 (see Appendix B for calculation details). In this model, the surface that separates the molten core from the outer H2-rich envelope is the binodal in the system MgSiO3-H2. The surface of the supercritical magma ocean is therefore as described by Young et al. (2024). This surface moves inward as the planet cools, resulting in a lower temperature and higher pressure for the phase boundary; the conditions at the binodal separating the magma ocean from the envelope are tied to the overall thermal state of the planet. Our fiducial model was chosen to represent an evolved planet with a relatively high pressure of 4 GPa for the binodal.
@@
Figure 9: Pressure as a function of radial position along an isentrope in the interior of our fiducial sub-Neptune with a fully miscible core. The envelope-core interface is at $R/R_{\rm c}=1$ where $R_{\rm c}$ is the radius of the interior exclusive of the outer hydrogen-rich envelope (the “core”). The planet total mass is $6M_{\oplus}$ and has a bulk H2 concentration of 2% by mass. The positions of the isobaric, isothermal phase diagrams in Figure 10 are shown for reference.